A gas present in a cylinder fitted with a frictionless piston expands against a constant pressure of 1 atm from a volume of 2 litre to a volume of 6 litre. In doing so, it absorbs 800J heat from surrounding. Determune increase in internal energy of process.

Grade:11

1 Answers

Arun
5 years ago

ΔH = ΔU + PΔV... (Definition of Enthalpy change)

At constant pressure (for isobaric process)

ΔH = q….

Hence, ΔU = q - PΔV (1st law of thermodynamics)

ΔU = 800 - 0.5*101325 *(0.006–0.002)

ΔU = -602.65 J

All the sign conventions were considered and pressure and volume were converted to their respective SI units.

1 atm pressure = 101325 Pa

1 liter volume = 0.001 cubic m.